# Inference of Dependency Knowledge Graph for Electronic Health Records

### v1

@@ 1.1 Our Contributions
The dynamic log-linear topic model, foundational to our work, has been adapted in several subsequent studies (Arora et al., 2018; Lu et al., 2023; Zhou et al., 2023). However, these adaptations have primarily focused on estimation, leaving a gap in uncertainty quantification. Undertaking inferential analysis on KGs presents three major challenges. For data generation, the model’s hierarchical structure entangles several time series and brings the data with a complex dependency structure. For estimation, two layers of non-linearity brought about by the logarithmic and division operators shown later in (2.4) further complicate the analysis. For data privacy, in many situations, we do not have access to patient-level data thus causing the estimation of covariance structure a problem.
Our work contributes to the field of statistical inference in low-rank models, a domain fraught with challenges. Performing statistical inference in such models is particularly difficult due to the intricate nature of low-rank structures and the complexity of their estimation processes. Recent advances have been made in linear low-rank models, including matrix completion (Foucart et al., 2017; Chen et al., 2019), principal component analysis (Xia, 2021; Xia et al., 2022) and low-rank matrix regression (Carpentier et al., 2019; Chernozhukov et al., 2023). However, extending these inferential techniques to non-linear pointwise mutual information under temporal dependent models remains a formidable challenge.
Our research makes several pivotal contributions. First, we enhance the understanding of the low-rank nature of the population PMI by deriving a shaper approximation rate using the KG embeddings. In specific, instead of assuming certain prior of the embeddings in the existing works (Arora et al., 2018; Lu et al., 2023), the embeddings in our model are deterministic vectors and thus make hypotheses on the cosine similarities between embeddings well-defined. The sharper rate is also crucial for valid inference. Second, we provide a method for uncertainty quantification of the estimator by demonstrating its asymptotic normality, addressing a significant gap where no previous valid inference method for non-linear low-rank models. The complexity of this analysis stems from the high non-linearity of the generative model and the dependencies inherent in longitudinal EHR datasets. Lastly, our study introduces an inference procedure for the KG using only summary-level EHR data. This approach is especially relevant when patient-level data sharing is not feasible, allowing for effective analysis that adheres to privacy constraints.
@@ 1.2 Paper Organization
The rest of the paper is organized as follows. Section 2 introduces the generative model of EHR data and proposes the method for KG inference. Section 3 provides the theoretical properties of the empirical PMI matrix and the entry-wise asymptotic normality of the low-rank estimator. Section 4 shows the simulation results and Section 5 applies the proposed method to construct a KG of four domains of EHR entities based on the empirical PMI matrix along with other summary statistics derived from EHR data of over $12$ million patients in the Veteran Affairs (VA) health system. The conclusion is given in Section 6. In addition, the proof of the theorems is developed in the supplementary.
@@ 2.1 Notation
We use $\boldsymbol{1}_{p}$ to represent a $p$-dimensional vector with all entries equal to $1$ and $\mathbf{I}_{p}$ to represent the $p\times p$ identity matrix. Let $\mathcal{O}^{p\times p}=\{\mathbf{R}\in\mathbb{R}^{p\times p}:\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}_{p}\}$ denote the set of all $p\times p$ orthonormal matrix. For a vector $\bm{a}\in\mathbb{R}^{p}$, we use $\|\bm{a}\|=\sqrt{\sum_{i=1}^{p}a_{i}^{2}}$ to represent its $l_{2}$ norm. For a matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$, we denote by $\mathbf{A}_{i,\cdot}=(A_{i1},\cdots,A_{im})$ (respectively $\mathbf{A}_{\cdot,i}$) its $i$th row (respectively column), $\mathbf{A}_{i,i^{\prime}}$ its $(i,i^{\prime})$th element, $\|\mathbf{A}\|_{\max}=\max_{i,j}|\mathbf{A}_{ij}|$ its matrix max norm, $\|\mathbf{A}\|_{2,\infty}=\max_{i}\|\mathbf{A}_{i,\cdot}\|$ its $2$-to-$\infty$ norm, and $\|\mathbf{A}\|$ its matrix operator norm. The Hadamard product of matrices $\mathbf{A}$ and $\mathbf{B}$ is denoted by $\mathbf{A}\circ\mathbf{B}$. If a symmetric matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ has eigen-decomposition $\mathbf{U}{\boldsymbol{\Lambda}}\mathbf{U}^{\top}$, where ${\boldsymbol{\Lambda}}=\text{diag}(\lambda_{1},\cdots,\lambda_{n})$, and $|\lambda_{1}|\geq|\lambda_{2}|\geq\cdots\geq|\lambda_{n}|$ (sort by magnitude), then we use $\lambda_{i}(\mathbf{A})=\lambda_{i}$ to represent the $i$th largest absolute eigenvalue of $\mathbf{A}$. For two sequences $\{x_{n}\}$ and $\{y_{n}\}$, we say $x_{n}=O(y_{n})$ or $x_{n}\lesssim y_{n}$ if there exists some constant $C>0$ such that $x_{n}\leq Cy_{n}$ for all $n$. We say $x_{n}=o(y_{n})$ if $\lim_{n\rightarrow\infty}x_{n}/y_{n}=0$, and $x_{n}=\Omega(y_{n})$ if $y_{n}=O(x_{n})$. Denote $[d]$ as the set $\{1,2,\cdots,d\}$. Suppose there are a total of $n$ patients and $d$ unique EHR entities. We let $\Phi(\cdot)$ be the cumulative distribution function of the standard normal distribution.
@@ 2.2 Generation of the Longitudinal EHR Feature Occurrences
In this section, we introduce the generative model for the observed longitudinal EHR data. We assume the entities in the EHR are generated from the dynamic log-linear topic model (Arora et al., 2016; Lu et al., 2023). Specifically, for the $i\in[n]$th patient, we observe the longitudinal EHR data $\{w_{i,1},\ldots,w_{i,t},\ldots,w_{i,T_{i}}\}$, where $T_{i}$ is the last observation time, and $w_{i,t}$ is the EHR entity observed at time $t$ and follows the following multinomial distribution
@@
where $p$ is the dimension of the embedding, $\alpha=1-(\log d)/p^{2}$ is a measure of the walking rate aligning with that in Lu et al. (2023), $\boldsymbol{r}_{i,t}$ is independent of $\{{\boldsymbol{c}}_{i,k}\}_{k=1}^{t-1}$; and the discourse vectors are independent across patients. This discourse process is stationary under such setting. We consider the sequence $\{{\boldsymbol{c}}_{i,t}\}_{t=1}^{T_{i}}$ follows an AR model different from the sphere random walk as in Arora et al. (2016) to incorporate the temporal variance heterogeneity of the discourse.
Denote the KG embedding matrix by $\mathbf{V}=(\mathbf{V}_{1},\ldots,\mathbf{V}_{d})^{\top}$. We see from (2.1) that the probability distribution over the entities is invariant to a constant shift of the rows of $\mathbf{V}$, i.e., for any constant vector $\bm{\mu}\in\mathbb{R}^{p}$, replacing $\mathbf{V}$ with $\mathbf{V}-\boldsymbol{1}_{d}\bm{\mu}^{\top}$ does not change the probability of occurrence. In order to have the identifiable embeddings, we assume with out loss of generality that $\mathbf{V}$ is centered, i.e., $\mathbf{V}^{\top}\bm{p}=\mathbf{0}$, where $\bm{p}=(p_{1},\cdots,p_{d})^{\top}\in\mathbb{R}^{d}$ with $p_{w}={\mathbb{E}}_{{\boldsymbol{c}}_{i,t}\sim N(0,\mathbf{I}_{p}/p)}[{\mathbb{P}}(w_{i,t}=w|{\boldsymbol{c}}_{i,t})]$ being the marginal occurrence probability of feature $w$ over the stationary discourse process. Thus the identifiability condition $\mathbf{V}^{\top}\bm{p}=0$ implies that the feature embeddings are centered at zero under to the distribution of marginal occurrence probability $\bm{p}$.
@@
Our approach, encapsulated in the hypothesis testing procedure (2.3), addresses this gap. By applying this hypothesis testing to all entity pairs $(w,w^{\prime})$ where $w\neq w^{\prime}$ within the set $[d]\times[d]$, we can construct a sparsely connected KG. This testing is equivalent to assessing the cosine similarity
@@ 2.3 Inferential Procedure through Low-Rank Approximation
$\mathbb{C}_{w,w^{\prime}}=\sum_{i=1}^{n}\mathbb{C}^{(i)}_{w,w^{\prime}},$
[…]
We then denote $\mathbb{C}^{(i)}=[\mathbb{C}^{(i)}_{w,w^{\prime}}]_{w,w^{\prime}\in[d]}$ as the $i$th patient’s co-occurrence matrix, $\mathbb{C}=[\mathbb{C}_{w,w^{\prime}}]_{w,w^{\prime}\in[d]}$ as the total co-occurrence matrix, $\mathbb{C}^{(i)}_{w}=\sum_{w^{\prime}}\mathbb{C}^{(i)}_{w,w^{\prime}}$ the individual marginal occurrence of $w$, and $\mathbb{C}_{w}=\sum_{w^{\prime}=1}^{d}\mathbb{C}_{w,w^{\prime}}$ the marginal occurrence of $w$, and finally, ¯$\overline{\mathbb{C}}^{(i)}=\sum_{w}\mathbb{C}^{(i)}_{w}$ and ¯$\overline{\mathbb{C}}=\sum_{w}\mathbb{C}_{w}$.
@@ 2.3.2 Variance Estimation without Patient-level Data
where $\widehat{\bm{p}}=(\widehat{p}_{1},\cdots,\widehat{p}_{d})^{\top}$, $T_{0}=Tq-q(q+1)/2$ and $T=\sum_{i=1}^{n}T_{i}/n$. To estimate the covariance matrices of $\mathbf{W}$, we only need to know parameters $n,T,q,$ and the estimate of $\{p_{w}\}_{w=1}^{d}$, which can be obtained from the summary-level co-occurrence matrix $\mathbb{C}$ by (2.6). With $\widehat{\text{Cov}}(\mathbf{W}_{w,\cdot},\mathbf{W}_{w^{\prime},\cdot})$ obtained, we can further estimate the variance of $\widetilde{\mathbb{PMI}}(w,w^{\prime})$ by (2.10) with $\widehat{\bm{\Sigma}}_{w,w^{\prime}}$ replaced by $\breve{\bm{\Sigma}}_{w,w^{\prime}}=\widehat{\text{Cov}}\big{(}\mathbf{W}_{w\cdot},\mathbf{W}_{w^{\prime}\cdot}\big{)}$ as
@@
and conduct the hypothesis testing based on the asymptotic normality of $\widetilde{\mathbb{PMI}}$. The KNowledge graph Inference and Testing (KNIT) procedure is then summarized in Algorithm 1.
@@
  Input: the summary-level co-occurrence matrix $\mathbb{C}\in\mathbb{R}^{d\times d}$, the window size $q$, the number of patients $n$, the rank of KG embeddings $r$, the average observation time $T=\sum_{i=1}^{n}T_{i}/n$, and the nominal FDR level $\alpha$.
@@
  Inference: Obtain the $p$-value $\alpha_{w,w^{\prime}}=\Phi\big{(}\widetilde{\mathbb{PMI}}(w,w^{\prime})/\sqrt{\widetilde{\rm{Var}}(\widetilde{\mathbb{PMI}}(w,w^{\prime}))}\big{)}$ for $1\leq w<w^{\prime}\leq d$. Order $\alpha_{w,w^{\prime}}$’s as $\alpha_{(1)}\leq\alpha_{(1)}\leq\ldots\leq\alpha_{(J)}$ with $J=d(d-1)/2$ and set $\alpha_{(0)}=0$. Let $j_{\max}=\max\{0\leq j\leq J:\alpha_{(j)}\leq\alpha j/J/(\log J+1)\}$.

### v2

@@ 1.1 Our Contributions
The dynamic log-linear topic model, foundational to our work, has been adapted in several studies (Arora et al., 2018; Lu et al., 2023; Zhou et al., 2023). However, these adaptations have primarily focused on estimation, leaving a gap in uncertainty quantification. Undertaking inferential analysis on KGs presents three major challenges. For data generation, the model’s hierarchical structure entangles several time series and brings the data with a complex dependency structure. For estimation, two layers of non-linearity brought about by the logarithmic and division operators shown later in (2.4) further complicate the analysis. For data privacy, in many situations, we do not have access to patient-level data thus causing the estimation of covariance structure a problem.
Our work contributes to the field of statistical inference in low-rank models, a domain fraught with challenges. Performing statistical inference in such models is particularly difficult due to the intricate nature of low-rank structures and the complexity of their estimation processes. Recent advances have been made in linear low-rank models, including matrix completion (Foucart et al., 2017; Chen et al., 2019), principal component analysis (Xia, 2021; Xia et al., 2022) and low-rank matrix regression (Carpentier et al., 2019; Chernozhukov et al., 2023). These work assume that the population matrix itself is low-rank. In contrast, our setting is fundamentally different - the population PMI matrix is not exactly low-rank but emerges from a generative model. We assume that the embedding dimension is low compared to the vocabulary size, which is widely used in both theoretical analysis and in reality where the embedding dimension is only several hundreds and the vocabulary size is more than tens of thousands. We prove that under our generative model, the population PMI matrix can be well approximated by a low-rank matrix, aligning with empirical observations in previous studies (Levy and Goldberg, 2014; Arora et al., 2016). We develop a novel inference framework that leverages a second-order approximation of the PMI matrix and employs advanced perturbation techniques, such as Davis-Kahan bounds, to establish theoretical guarantees. Unlike existing approaches, our method extends inferential techniques to non-linear PMI estimation under temporal dependent generative models, where the ground truth matrix is derived from the generative process rather than assumed to be low-rank.
Our research makes several pivotal contributions. First, we enhance the understanding of the near low-rank nature of the population PMI by deriving a shaper approximation rate using KG embeddings. While prior works (Arora et al., 2016; Lu et al., 2023) have explored the approximate low-rank structure of the population PMI matrix, existing analyses primarily focus on two aspects: the empirical PMI estimator without a low-rank decomposition and the estimation consistency, without extending to statistical inference. The key challenge preventing inference lies in the variance-bias decomposition. Previous approaches relied on a first-order Taylor expansion to approximate the PMI matrix with an inner product of the embeddings. However, the bias term from the first-order expansion is larger than the standard deviation of the estimator, making inference infeasible. A key contribution of our work is the discovery that a second-order Taylor expansion still results in a low-rank structure but significantly reduces the bias, making inference possible. Moreover, instead of assuming certain prior of the embeddings in the existing works (Arora et al., 2018; Lu et al., 2023), the embeddings in our model are deterministic vectors and thus make hypotheses on the cosine similarities between embeddings well-defined. The sharper rate is also crucial for valid inference.
Second, we provide a method for uncertainty quantification of the estimator by demonstrating its asymptotic normality, addressing a significant gap where no previous valid inference method for non-linear low-rank models with complex data dependency. The complexity of this analysis stems from the high non-linearity of the generative model and the dependencies inherent in longitudinal EHR datasets. The entries of the PMI matrix are non-linear transforms of the co-occurrence counts involving the softmax and log transformations. In the Supplementary S.1 and S.2.1, we develop a series of novel concentration inequalities for the nonlinear transformation of samples from a geometrically ergodic Markov chain. This inequality plays a crucial role in establishing the asymptotic normality of our PMI estimator. As the empirical PMI matrix is obtained from the occurrence data generated by a latent Markovian process, we cannot apply the standard proof technique to show the asymptotic normality. In Supplementary S.2.2, we propose a new theoretical analysis to decouple the temporal-dependence of the discourse vectors and show the asymptotic normality as long as quantifying the scale of their asymptotic variances.
Lastly, our study introduces an inference procedure for the KG using only summary-level EHR data. This approach has significant practical value in healthcare settings where sharing patient-level data is restricted due to privacy regulations. By enabling rigorous statistical inference while preserving privacy, our method provides both a theoretically sound and practically implementable solution for analyzing multi-source EHR data.
In summary, our work goes beyond existing literature by (i) developing a second-order approximation that enables statistical inference on PMI matrices, (ii) addressing variance estimation and asymptotic normality in the presence of non-linearity and time dependencies, and (iii) proposing a privacy-preserving inference framework for knowledge graph construction. These contributions make our findings both theoretically novel and practically relevant.
@@ 1.2 Paper Organization
The rest of the paper is organized as follows. Section 2 introduces the generative model of EHR data and proposes the method for KG inference. Section 3 provides the theoretical properties of the empirical PMI matrix and the entry-wise asymptotic normality of the low-rank estimator. Section 4 shows the simulation results and Section 5 applies the proposed method to construct a KG of four domains of EHR entities based on the empirical PMI matrix along with other summary statistics derived from EHR data of over $12$ million patients in the Veteran Affairs (VA) health system. The conclusion is given in Section 6. In addition, the proof of the theorems is developed in the supplementary, and the implementation code is available at https://github.com/junwei-lu/WordVec_Inference.
@@ 2.1 Notation
We use $\boldsymbol{1}_{p}$ to represent a $p$-dimensional vector with all entries equal to $1$ and $\mathbf{I}_{p}$ to represent the $p\times p$ identity matrix. Let $\mathcal{O}^{p\times p}=\{\mathbf{R}\in\mathbb{R}^{p\times p}:\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}_{p}\}$ denote the set of all $p\times p$ orthonormal matrix. For a vector $\bm{a}\in\mathbb{R}^{p}$, we use $\|\bm{a}\|=\sqrt{\sum_{i=1}^{p}a_{i}^{2}}$ to represent its $l_{2}$ norm. For a matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$, we denote by $\mathbf{A}_{i,\cdot}=(A_{i1},\cdots,A_{im})$ (respectively $\mathbf{A}_{\cdot,i}$) its $i$th row (respectively column), $\mathbf{A}_{i,i^{\prime}}$ its $(i,i^{\prime})$th element, $\|\mathbf{A}\|_{\max}=\max_{i,j}|\mathbf{A}_{ij}|$ its matrix max norm, $\|\mathbf{A}\|_{2,\infty}=\max_{i}\|\mathbf{A}_{i,\cdot}\|$ its $2$-to-$\infty$ norm, and $\|\mathbf{A}\|$ its matrix operator norm. The Hadamard product of matrices $\mathbf{A}$ and $\mathbf{B}$ is denoted by $\mathbf{A}\circ\mathbf{B}$. If a symmetric matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ has eigen-decomposition $\mathbf{U}{\boldsymbol{\Lambda}}\mathbf{U}^{\top}$, where ${\boldsymbol{\Lambda}}=\text{diag}(\lambda_{1},\cdots,\lambda_{n})$, and $|\lambda_{1}|\geq|\lambda_{2}|\geq\cdots\geq|\lambda_{n}|$ (sort by magnitude), then we use $\lambda_{i}(\mathbf{A})=\lambda_{i}$ to represent the $i$th largest absolute eigenvalue of $\mathbf{A}$. For two sequences $\{x_{n}\}$ and $\{y_{n}\}$, we say $x_{n}=O(y_{n})$ or $x_{n}\lesssim y_{n}$ if there exists some constant $C>0$ such that $x_{n}\leq Cy_{n}$ for all $n$. We say $x_{n}=o(y_{n})$ if $\lim_{n\rightarrow\infty}x_{n}/y_{n}=0$, and $x_{n}=\Omega(y_{n})$ if $y_{n}=O(x_{n})$. Denote $[d]$ as the set $\{1,2,\cdots,d\}$. We let $\Phi(\cdot)$ be the cumulative distribution function of the standard normal distribution.
@@ 2.2 Generation of the Longitudinal EHR Feature Occurrences
In this section, we introduce the generative model for the observed longitudinal EHR data. Suppose that there are a total of $n$ patients and $d$ unique EHR entities. We use the set $d]=\{1,2,\ldots,d\}$ to represent the vocabulary of all possible EHR entities (e.g., diagnosis codes, medications, procedures) observed across the patient cohort. We assume that the entities in the EHR are generated from the dynamic log-linear topic model (Arora et al., 2016; Lu et al., 2023). Specifically, for the $i\in[n]$th patient, we observe the longitudinal EHR data as the sequence of EHR entities $\{w_{i,1},\ldots,w_{i,t},\ldots,w_{i,T_{i}}\}$, where $T_{i}$ is the last observation time, and $w_{i,t}$ denotes a single EHR entity (or code) observed at time $t$. An entity $w_{i,t}\in[d]$ could represent a diagnosis code, such as an ICD code, a medication code, or any other discrete piece of information recorded in the patient’s record at time $t$. For example, if $w_{i,t}$ is the diagnosis code of Hypertension, it might represent that the patient was diagnosed with “Hypertension” at time $t$. Similarly, if $w_{i,t}$ corresponds to a procedure code for “Blood pressure measuremen”, it indicates that the patient underwent a blood pressure measurement at time $t$. These discrete entities provide granular information about the patient’s medical history.
We assume that $w_{i,t}$ follows the following multinomial distribution
@@
where $p$ is the dimension of the embedding, $\alpha=1-(\log d)/p^{2}$ is a measure of the walking rate aligning with that in Lu et al. (2023), $\boldsymbol{r}_{i,t}$ is independent of $\{{\boldsymbol{c}}_{i,k}\}_{k=1}^{t-1}$; and the discourse vectors are independent across patients. This discourse process is stationary in such a setting. We consider that the sequence $\{{\boldsymbol{c}}_{i,t}\}_{t=1}^{T_{i}}$ follows an AR model different from the random walk of the sphere as in Arora et al. (2016) to incorporate the heterogeneity of the temporal variance of the discourse. Although this assumption facilitates theoretical analysis, our method is empirically robust to variations in the distribution of ${\boldsymbol{c}}_{i,k}$, as demonstrated in Figure 2.
Remark 2.1 (Suitability of the Model for EHR Data).
Although the model in (2.1) was originally developed for text generation (Arora et al., 2016), it is particularly well-suited for modeling EHR data. Like text data, EHR data consist of sequences of discrete events or entities, such as diagnoses, procedures, or medications, collected over time. The dynamic log-linear topic model effectively captures the temporal dependencies and latent structures inherent in these sequences, making it a natural choice for analyzing longitudinal EHR data.
A key strength of this model is the inclusion of a latent discourse vector ${\boldsymbol{c}}_{i,t}$, which evolves over time to reflect changes in a patient’s medical status. This feature is particularly important in healthcare, where disease progression and treatment responses unfold over time. By modeling ${\boldsymbol{c}}_{i,t}$ as an autoregressive process, the framework accommodates heterogeneous temporal variances in patient records, allowing for more accurate modeling of real-world medical trajectories.
Finally, the integration of knowledge graph embeddings $\mathbf{V}_{w}$ enhances the interpretability of the model. These embeddings enable the inference of relationships between medical entities, facilitating the discovery of hidden connections among diagnoses, treatments, and outcomes. This makes the model especially useful for extracting clinically meaningful insights from large-scale EHR datasets.
Denote the KG embedding matrix by $\mathbf{V}=(\mathbf{V}_{1},\ldots,\mathbf{V}_{d})^{\top}$. We see from (2.1) that the probability distribution over the entities is invariant to a constant shift of the rows of $\mathbf{V}$, i.e., for any constant vector $\bm{\mu}\in\mathbb{R}^{p}$, replacing $\mathbf{V}$ with $\mathbf{V}-\boldsymbol{1}_{d}\bm{\mu}^{\top}$ does not change the probability of occurrence. In order to have identifiable embeddings, we assume without loss of generality that $\mathbf{V}$ is centered, i.e., $\mathbf{V}^{\top}\bm{p}=\mathbf{0}$, where $\bm{p}=(p_{1},\cdots,p_{d})^{\top}\in\mathbb{R}^{d}$ with $p_{w}={\mathbb{E}}_{{\boldsymbol{c}}_{i,t}\sim N(0,\mathbf{I}_{p}/p)}[{\mathbb{P}}(w_{i,t}=w|{\boldsymbol{c}}_{i,t})]$ being the marginal occurrence probability of feature $w$ over the stationary discourse process. Thus, the identifiability condition $\mathbf{V}^{\top}\bm{p}=0$ implies that the feature embeddings are centered at zero under the marginal occurrence probability $\bm{p}$.
@@
Our approach, encapsulated in the hypothesis testing procedure (2.3), addresses this gap. By applying this hypothesis testing to all entity pairs $(w,w^{\prime})$ where $w\neq w^{\prime}$ within the set $d]\times[d]$, we can construct a sparsely connected KG. This testing is equivalent to assessing the cosine similarity
@@ 2.3 Inferential Procedure through Low-Rank Approximation
$\mathbb{C}_{w,w^{\prime}}=\sum_{i=1}^{n}\mathbb{C}^{(i)}_{w,w^{\prime}},\text{ where }$
$\mathbb{C}^{(i)}_{w,w^{\prime}}=|\{(t,s)\in[T_{i}]\times[T_{i}]:0<|t-s|\leq q,\min\{t,s\}\leq T_{i}-q\text{ and }w_{i,t}=w,w_{i,s}=w^{\prime}\}|.$
We then denote $\mathbb{C}^{(i)}=[\mathbb{C}^{(i)}_{w,w^{\prime}}]_{w,w^{\prime}\in[d]}$ as the $i$th patient’s co-occurrence matrix, $\mathbb{C}=[\mathbb{C}_{w,w^{\prime}}]_{w,w^{\prime}\in[d]}$ as the total co-occurrence matrix, $\mathbb{C}^{(i)}_{w}=\sum_{w^{\prime}}\mathbb{C}^{(i)}_{w,w^{\prime}}$ the individual marginal occurrence of $w$, and $\mathbb{C}_{w}=\sum_{w^{\prime}=1}^{d}\mathbb{C}_{w,w^{\prime}}$ the marginal occurrence of $w$, and finally, $\overline{\mathbb{C}}^{(i)}=\sum_{w}\mathbb{C}^{(i)}_{w}$ and $\overline{\mathbb{C}}=\sum_{w}\mathbb{C}_{w}$.
@@ 2.3.2 Variance Estimation without Patient-level Data
where $\widehat{\bm{p}}=(\widehat{p}_{1},\cdots,\widehat{p}_{d})^{\top}$, $T_{0}=Tq-q^{2}$ and $T=\sum_{i=1}^{n}T_{i}/n$. To estimate the covariance matrices of $\mathbf{W}$, we only need to know parameters $n,T,q,$ and the estimate of $\{p_{w}\}_{w=1}^{d}$, which can be obtained from the summary-level co-occurrence matrix $\mathbb{C}$ by (2.6). With $\widehat{\text{Cov}}(\mathbf{W}_{w,\cdot},\mathbf{W}_{w^{\prime},\cdot})$ obtained, we can further estimate the variance of $\widetilde{\mathbb{PMI}}(w,w^{\prime})$ by (2.10) with $\widehat{\bm{\Sigma}}_{w,w^{\prime}}$ replaced by $\breve{\bm{\Sigma}}_{w,w^{\prime}}=\widehat{\text{Cov}}\big(\mathbf{W}_{w\cdot},\mathbf{W}_{w^{\prime}\cdot}\big)$ as
@@
and conduct the hypothesis testing based on the asymptotic normality of $\widetilde{\mathbb{PMI}}$. The KNowledge graph Inference and Testing (KNIT) procedure is then summarized in Algorithm 1, where the rank of KG embeddings $p$ is estimated by Algorithm 2.
@@
Input: the summary-level co-occurrence matrix $\mathbb{C}\in\mathbb{R}^{d\times d}$, the window size $q$, the number of patients $n$, the rank of KG embeddings $p$, the average observation time $T=\sum_{i=1}^{n}T_{i}/n$, and the nominal FDR level $\alpha$.
@@
Inference: Obtain the $p$-value $\alpha_{w,w^{\prime}}=\Phi\big(\widetilde{\mathbb{PMI}}(w,w^{\prime})/\sqrt{\widetilde{\rm{Var}}(\widetilde{\mathbb{PMI}}(w,w^{\prime}))}\big)$ for $1\leq w<w^{\prime}\leq d$. Order $\alpha_{w,w^{\prime}}$’s as $\alpha_{(1)}\leq\alpha_{(1)}\leq\ldots\leq\alpha_{(J)}$ with J=d(d1)/2J=d(d-1)/2 and set $\alpha_{(0)}=0$. Let $j_{\max}=\max\{0\leq j\leq J:\alpha_{(j)}\leq\alpha j/J/(\log J+1)\}$.
@@
Algorithm 2 Estimating the rank of knowledge graph embeddings
Input: Eigenvalues of $\widehat{\mathbb{PMI}}$ ($\widehat{\lambda}_{1},\ldots,\widehat{\lambda}_{d}$), thresholding parameter $\eta_{0}={8d^{3}p\log^{2}d}/{\sqrt{nT}}$.
Estimation: Compute $r=\max\{k\in[d]:\widehat{\lambda}_{k}\geq\eta_{0}\}$.
Output: The estimated rank of knowledge graph embeddings, $r$.
The following proposition guarantees that Algorithm 2 recovers the true rank of KG embeddings with high probability.
Proposition 2.1.
Under the condition that ${d^{3}p\log^{2}d}/{\sqrt{nT}}\ll\kappa^{2}\xi^{2}$, with probability at least $1-\exp(-\Omega(\log^{2}d))$, we have that $r=p$.
The proof follows directly from the proof of Proposition S.23 in the Supplementary Material.